To build the garden of your dreams you 10 ft3 of soil containing 17% clay. you have two types of soil you can combine to achieve this: soil with 35% clay and soil with 10% clay. how much of each soil should you use?
Answer by stanbon(75887)

To build the garden of your dreams you need 10 ft^3 of soil containing 17% clay. you have two types of soil you can combine to achieve this: soil with 35% clay and soil with 10% clay. how much of each soil should you use?
----
Equation:
clay + clay = clay
0.35x + 0.10(10-x) = 0.17*10
---
Multiply thru by 100 to get:
35x + 10*10 - 10x = 17*10
25x = 7*10
---
x = (2/5)7
x = 14/5 = 2.8 ft^3 (amt. of 35%-clay soil needed)
-----
10-2.8 = 7.2 ft^3 (amt of 10%-clay soil needed)
=========================================
cheers,
Stan H.